OpenStudy (anonymous):

The altitude of a triangle is 5 m less than its base. The area of the triangle is 80 m^2. Find the base. ( use quadratic equation form...) so far I got this: x^2-5x=80 the only problem is, it's a triangle, so i have to also add the 1/2 part but how do I do that?

OpenStudy (lgbasallote):

\[\frac{x^2 - 5x}{2} = 80\]multipkly both sides by 2 \[x^2 - 5x = 160\] that answer your question?

OpenStudy (anonymous):

ahhhhhh NICEE LOL THANKS :D

OpenStudy (lgbasallote):

it's like that because it's bh/2 b = x h = x - 5

OpenStudy (anonymous):

I M ALSO NOT GETTING THIS

OpenStudy (anonymous):

THANKS FOR THE HELP @lgbasallote

OpenStudy (lgbasallote):

just doing my avatar duties :D